# An unsupervised posterior sampling framework for multi-purpose seismic data recovery

### v1

@@ Abstract
Seismic data restoration is a fundamental task in seismic exploration, yet remains challenging under complex and unknown degradations. Traditional model-driven or task-specific learning methods often require retraining for each degradation type and fail to generalize effectively to unseen field data. In this work, we introduce an unsupervised Posterior Sampling Framework (PSF) built upon Score-based Generative Models (SGMs) for unified seismic data restoration. PSF leverages a pre-trained unconditional SGMs as a seismic-aware generative prior and derives a generalized conditional score function linked to the forward operator of each inverse problem. This enables posterior sampling across different seismic restoration tasks without retraining or supervision. Additionally, an adaptive noise-level estimation mechanism is incorporated to dynamically regulate the noise suppression strength during sampling, enhancing flexibility under varying signal-to-noise ratios and degradation conditions. Extensive experiments on seismic denoising, interpolation, compressed sensing, and deconvolution demonstrate that PSF delivers high-quality samples and exhibits robust generalization to out-of-distribution data. These results highlight the potential of SGMs as a universal prior for seismic inverse problems and establish PSF as a flexible framework for unsupervised posterior inference across diverse degradation scenarios.
@@ 2 Introduction
These limitations of both traditional model-driven methods and deep learning approaches highlight the need for more flexible and generalizable frameworks capable of handling diverse seismic inverse problems without requiring task-specific retraining or strong prior assumptions. Recently, Score-based Generative Models (SGMs),also known as diffusion models, have emerged as a powerful class of generative priors capable of modeling complex, high-dimensional data distributions through score functions [21, 22, 23]. SGMs have been increasingly adopted in geophysics for generative modeling of various types of seismic data ([24, 25, 26]). SGMs have demonstrated remarkable performance across a wide range of seismic inverse problems, including denoising [27, 28], interpolation  [29, 30, 31, 32, 33], resolution enhancement [34, 35], impedance inversion [36], and seismic imaging [37]. Despite these advances, existing diffusion-based approaches still face several fundamental limitations. Most current methods remain task-specific, relying on paired training data and costly retraining when adapting to new degradation types [29, 34, 37, 32, 33], which hinders their ability to generalize online or be deployed in real-world seismic settings. Furthermore, they typically design task-dependent conditional score functions rather than a unified posterior sampling framework applicable to diverse inverse problems [28, 31]. As a result, once the forward operator changes, the entire sampling scheme must often be redesigned. In addition, few existing approaches provide a theoretical or practical mechanism for handling inverse problems under arbitrary or unknown noise levels, which constrains their robustness in field data scenarios [29, 30, 34, 36, 32, 33, 37].
@@
Despite recent progress, how to effectively extend a pre-trained unconditional SGMs to diverse seismic inverse problems remains largely unexplored. The key challenges lie in how to incorporate different forward operators associated with seismic degradations, such as noise injection, subsampling, convolution, without retraining, and how to adaptively regulate the level of noise suppression during posterior sampling to handle data with varying signal-to-noise ratios (SNRs). To address these challenges, this paper proposes an unsupervised Posterior Sampling Framework (PSF) based on pre-trained Score-based Generative Models (SGMs) for seismic data restoration. PSF treats the pre-trained unconditional SGMs as a universal generative prior and derives a generalized conditional score function corresponding to arbitrary forward operators. This allows posterior sampling for different inverse problems without model retraining. Furthermore, PSF integrates an automatic noise-level estimation strategy that adjusts the sampling process adaptively, thereby controlling the strength of noise suppression according to the estimated data noise. By bridging unconditional SGMs with task-specific forward operators, PSF provides a unified and flexible framework.
@@ 3.2.2 Posterior sampling using Langevin dynamics and conditional score function
In Sections 3.2.3 and 3.3, the forward operator $\mathbf{G}$ in Equation 8 is introduced, and its low-memory SVD decomposition is presented. Furthermore, in Section 3.4, an automatic noise-level estimation algorithm is provided to estimate the parameter $\sigma_{\mathbf{y}}$ in Equation 9.
@@
14:   $\mathbf{d}_{t}\Leftarrow\mathbf{\Sigma}^{T}\cdot\left|\sigma_{\mathbf{y}}^{2}\mathbf{I}-\sigma_{i}^{2}\mathbf{\Sigma\Sigma}^{T}\right|^{\dagger}\cdot\left(\mathbf{U}^{T}\mathbf{y}-\mathbf{\Sigma V}^{T}\mathbf{x}_{t-1}\right)+\left.\left(\mathbf{V}^{T}\cdot{{\bf{s}}_{{{\bf{\theta}}^{*}}}}\left(\mathbf{x}_{t-1},\sigma_{i}\right)\right)\right|_{\not{>}}$ $\triangleright$ ${\bf{s}}_{{{\bf{\theta}}^{*}}}$} is known
@@ 3.3 Memory Efficient SVD
Methods
PSF
SGMPS-RNA
[28]
SGMPS-SDR
[31]
SNIPS
[38]
Multi-purpose
posterior sampling
$\checkmark$ $\times$ $\times$ $\checkmark$
@@ Methods
SNIPS (NCSNv2)
[38]
$128\times 128$ 86.07s 6.05it/s
@@
SNIPS (NCSNv2)
[38]
$256\times 256$ Out of memory
@@ 5 Discussion
PSF is a flexible, unsupervised posterior sampling framework. As shown in Figure 12, we demonstrate seismic data recovery under varying signal-to-noise ratios, different missing-trace patterns, diverse compression ratios, and convolutions with wavelets of different dominant frequencies. By automatically estimating $\sigma_{y}$ and adjusting the operator $\mathbf{G}$, PSF can handle seismic data with unknown noise levels and unknown degradation processes. Beyond the four tasks presented in this work, PSF can be extended to additional applications, such as super-resolution and deblurring, by modifying $\mathbf{G}$ (see Figure 13). For more complex nonlinear inverse problems, $\mathbf{G}$ can first be linearized before applying PSF for reconstruction. Since PSF requires knowledge of the forward operator $\mathbf{G}$, new posterior sampling algorithms are needed when $\mathbf{G}$ is unknown, such as in blind deconvolution tasks with unknown wavelets. Moreover, because PSF posterior sampling involves hundreds of function evaluations, developing accelerated sampling strategies will be essential. These directions represent promising avenues for future research.
Figure 12: Posterior sampling under different signal-to-noise ratios and different degradation operators
Figure 13: Posterior sampling examples for more tasks

### v2

# Unsupervised Posterior Sampling for Seismic Data Recovery via Score-Based Generative Priors

@@ Abstract
Seismic data restoration is a fundamental task in seismic exploration, yet remains challenging under complex and unknown degradations. Traditional model-driven or task-specific learning methods often require retraining for each degradation type and fail to generalize effectively to unseen field data. In this work, we introduce an unsupervised Posterior Sampling Framework (PSF) built upon Score-based Generative Models (SGMs) for unified seismic data restoration. PSF leverages the pre-trained unconditional SGMs as a seismic-aware generative prior and derives a generalized conditional score function associated with the forward operator of each inverse problem. This enables posterior sampling across different seismic restoration tasks without retraining or supervision. Additionally, an adaptive noise-level estimation mechanism is incorporated to dynamically regulate the noise suppression strength during sampling, enhancing flexibility under varying signal-to-noise ratios and degradation conditions. Extensive experiments on seismic denoising, interpolation, compressed sensing, and deconvolution demonstrate that PSF delivers high-quality samples and exhibits robust generalization to out-of-distribution data. These results highlight the potential of SGMs as a universal prior for seismic inverse problems and establish PSF as a flexible framework for unsupervised posterior inference across diverse degradation scenarios.
@@ 1 Introduction
Seismic data recovery can be viewed as an inverse problem. From a Bayesian perspective, recovering the useful signal corresponds to drawing samples from the posterior distribution, which provides a distribution of plausible solutions and naturally facilitates uncertainty quantification. In this context, posterior sampling offers a principled way to incorporate prior knowledge, while deep generative models provide flexible priors that bridge classical Bayesian inference with modern data-driven approaches. Conventional Bayesian seismic inversion relies on sampling techniques, such as Markov chain Monte Carlo, Gibbs sampling, and Langevin-type algorithms, to perform posterior inference and characterize uncertainty [13, 44, 26]. These approaches typically operate without leveraging deep generative priors, relying instead on analytic or hand-crafted prior formulations. Recent advances in deep generative modeling introduce data-driven priors that offer more expressive representations of seismic structures, providing an additional pathway for posterior sampling in seismic inversion. Early studies employing generative adversarial networks (GAN)[21] demonstrated that GAN-based priors enable ultrahigh-resolution or geologically realistic inversion results by constraining the solution space to a learned latent manifold [30, 55, 15]. These latent-variable priors integrate naturally with Bayesian posterior sampling, reducing the dimensionality of the model space while preserving structural realism. Variational autoencoders (VAE)[29] have likewise been used to capture geological variability and enable efficient posterior exploration through low-dimensional latent representations[34], with recent studies demonstrating that VAE-based dimensionality reduction can further improve uncertainty quantification in nonlinear geophysical inversion [32]. Similarly, normalizing-flow-based methods[11] have been adopted for seismic tomography and amplitude-versus-offset inversion[3, 61], leveraging invertible architectures to flexibly approximate complex, multimodal posterior distributions with significantly reduced computational cost. Collectively, these VAE-, GAN-, and flow-based approaches replace hand-crafted priors with learned generative priors, enabling scalable posterior inference and improved uncertainty quantification in seismic inversion.
More recently, score-based generative models (SGMs) [46, 24] have emerged as a highly flexible and expressive class of generative priors. By directly estimating the gradients of complex, high-dimensional data distributions, SGMs enable principled posterior sampling through stochastic differential equations [47], providing a unified and broadly applicable framework for Bayesian inversion across diverse seismic inverse problems. In geophysics, SGMs have demonstrated impressive performance across denoising [41, 39], interpolation [33, 54, 37, 53, 45], resolution enhancement [59, 56], impedance inversion [7], and seismic imaging [4]. Their strength lies in representing highly flexible priors that can be adapted to multiple inverse problems by incorporating forward operators during posterior sampling, without requiring task-specific retraining.
Nevertheless, existing SGM-based methods remain limited in flexibility and robustness. Many are trained for specific degradation types or tasks, necessitating retraining when the forward operator changes [33, 59, 4]. Conditional score functions are often designed for specific tasks, and few methods provide systematic strategies for handling unknown or variable noise levels, which is critical for field seismic data applications [33, 54, 7, 53, 45, 4]. While pre-training and fine-tuning paradigms can adapt generative models to new tasks, they still require supervised training with paired data or self-supervised learning during fine-tuning, and thus do not constitute fully unsupervised posterior sampling. Moreover, during the sampling stage, these generative models cannot simultaneously perform denoising and inversion[10].
@@
To address these challenges, we propose an unsupervised Posterior Sampling Framework (PSF) for seismic data restoration. PSF treats a pre-trained unconditional SGM as a universal generative prior and constructs a generalized posterior sampling scheme compatible with arbitrary forward operators. This allows flexible restoration of seismic signals under diverse degradation scenarios without retraining. Moreover, PSF integrates an automatic noise-level estimation mechanism to adaptively control the strength of noise suppression during sampling, ensuring robust performance across datasets with varying signal-to-noise ratios. By unifying unconditional SGMs with seismic-aware forward operators, PSF provides a theoretically grounded, flexible, and generalizable framework for seismic inverse problems.
@@ 2.2.2 Posterior sampling using Langevin dynamics and conditional score function
In Sections 2.2.3 and 2.3, the forward operator $\mathbf{G}$ in Equation 8 is introduced, and its low-memory SVD decomposition is presented. Furthermore, in Section 2.4, an automatic noise-level estimation algorithm is provided to estimate the parameter $\sigma_{\mathbf{y}}$ in Equation 9.
@@
14:   $\mathbf{d}_{t}\Leftarrow\mathbf{\Sigma}^{T}\cdot\left|\sigma_{\mathbf{y}}^{2}\mathbf{I}-\sigma_{i}^{2}\mathbf{\Sigma\Sigma}^{T}\right|^{\dagger}\cdot\left(\mathbf{U}^{T}\mathbf{y}-\mathbf{\Sigma V}^{T}\mathbf{x}_{t-1}\right)+\left.\left(\mathbf{V}^{T}\cdot{{\bf{s}}_{{{\bf{\theta}}^{*}}}}\left(\mathbf{x}_{t-1},\sigma_{i}\right)\right)\right|_{\leq}$ $\triangleright$ ${\bf{s}}_{{{\bf{\theta}}^{*}}}$} is known
@@ 2.3 Memory Efficient SVD
Methods PSF (proposed) SGMPS-RNA [39] SGMPS-SDR [37] SNIPS [28]
Multi-purpose posterior sampling $\checkmark$ $\times$ $\times$ $\checkmark$
@@
SNIPS (NCSNv2) [28] $128\times 128$ 86.07s 6.05it/s
@@
SNIPS (NCSNv2) [28] $256\times 256$ Out of memory
@@ 4 Discussion
PSF is a flexible, unsupervised posterior sampling framework. As shown in Figure 12, we demonstrate seismic data recovery under varying signal-to-noise ratios, different missing-trace patterns, diverse compression ratios, and convolutions with wavelets of different dominant frequencies. By automatically estimating $\sigma_{y}$ and adjusting the operator $\mathbf{G}$, PSF can handle seismic data with unknown noise levels and unknown degradation processes. For more complex nonlinear inverse problems, $\mathbf{G}$ can first be linearized before applying PSF for reconstruction. Since PSF requires knowledge of the forward operator $\mathbf{G}$, new posterior sampling algorithms are needed when $\mathbf{G}$ is unknown, such as in blind deconvolution tasks with unknown wavelets. Moreover, because PSF posterior sampling involves hundreds of function evaluations, developing accelerated sampling strategies will be essential. These directions represent promising avenues for future research.
Figure 12: Posterior sampling under different signal-to-noise ratios and different degradation operators, the ground truch for seismic data and reflection coefficients is shown in Figure 13.
Figure 13: the ground truch for seismic data and reflection coefficients in Figure 12